to 2 decimai places before multiplying by one and one-haif to determine the overtime rate) 2. If the third S. I the third dedimal place is less than 5, simply drop the third decimal place decimal place is 5 or more, round to the next higher cent Exampies: Monthly rate $1,827 Weekly rate ($1,827 x 12)/52 $421.615 rounded to $421.62 Hourly rate $421.62/40 $10.540 rounded to $10.54 Ο.Τ. rate $10.54 x 1.5 $15.81 Also, use the minimum hourty wage of $7.25 in solving these problems and all that follow Example 2-2 James Rudolph, a waiter, earns a weekly wage rate paid to Rudolph ($100.00 40 of s100.00 for a 40-hour workweek and receives tips of $100.00 eacn week. Even though the houry hours- $2.so) exceeds the minimum woge rate ofr $2 13, the employer must pay Rudoipn an extra seo.co so exceeds the minimum that the tips received plus the wages paid equal the minimum wage of $7 25 per hour 57.25 × 40 Weekly wage Tips received Extra weekly pay *S100.00 + S90.00 $190.00; $190.00 $290.00 (100.00) * (100.00) $90.00 40 hours $4.75/hour where he receives a weekly wage of $75 plus tips for a 40-hour workweek Garros weeky tips usually range from $300 to $350 Bert Garro is a weiter at La Bron H Round your answers to the nearest cent a. Under the Fair Labor Standerds Act, the minimum amount of wages that House, Garro must receive for a 40-hour workwesk is 90 X b. Since La Bron House is in viotation of the FLSA, minimum tipped wages.) the additionai amount it should pay Garro each week to meet the minmum wage requirement for a spped empioyee is ( Hint Employer must an additional amount to bring current gross of $75 up to -1301 x Previous Next All work saved t for Grading 9:26 ▼ ■
Answer #1
a
Minimum amount of wages = 2.13*40 = $85.20
b
Additional amount = 85.20-75 = $10.20